# Dept. of Informatics King’s College London London, UK

### v1

@@ 3.2 Previous Work
In the last 20 years, the importance of research on deceptive AI has been emphasised in the AI community during a series of events, namely the 2015 AAAI Fall Symposium on Deceptive and Counter-Deceptive Machines [^2] See the proceedings at https://aaai.org/Press/Reports/Symposia/Fall/fs-15-03.php., and the two International Workshops on Deceptive AI [^3] See the joint proceedings in Sarkadi et al. (2021b). co-located with ECAI2020 and IJCAI2021. Having gone through the socio-cognitive components that need to be considered when modelling deceptive AI, we now describe the work at the intersection of deception and AI, grouping them into six categories to better understand what computational aspects and properties deception they address.
@@ 3.2.1 Deception in Agent Societies
Recent findings back up this perspective by showing how humans and artificial agents influence each other’s attitudes and behaviour. For example, Mell et al. (2018) focus on how the negotiation strategies of artificial agents determine humans to endorse artificial deception. Apparently and counter-intuitively, humans also tend to be more cooperative with deceptive machines than with friendly ones, according to Ishowo-Oloko et al. (2019). Another type of study, presented by Dras et al. (2010), looks at the ways in which deceptive language can be generated and detected in virtual systems.
@@ 3.2.2 Logical Aspects of Agent Deception
Works in the areas of Logic and Philosophy have strongly influenced the current state of the art approaches in deceptive AI. Most known are the works in knowledge representation, which have laid the foundations to formalise deceptive communication. These appear in Sakama et al. (2010), in Sakama (2015), and in Sakama et al. (2015), where the authors present the logical distinctions between different types of dishonest linguistic behaviour, and later introducing a model of causality for them Sakama (2021);Bonnet et al. (2021) sketch a logical theory of belief manipulation; Van Ditmarsch (2014) introduce a dynamic logic of lying that considers lies about factual propositions as well as lies about the beliefs of others; in Uckelman (2011) where the author uses the medieval concept of dubitatio to study deceptive agents; Jones (2015) proposes a formal model of self-deception that is consistent given certain epistemic conditions imposed on the agent that engages in self-deception; and Smith et al. (2016) describe a logic-based account of analysing conjuring tricks as form of deception, and identify substitutable elements and stable occlusion as crucial components to construct perceived impossibilities.
@@ 3.2.3 Strategies for Agent Deception
Some of the work in this area explores issues such as using heuristics to cause target agents to execute a plan that will achieve the deceiver’s desired goal Christian and Young (2004); finding deceptive strategies using path-planning Masters and Sardina (2017); Masters et al. (2021a); Benke et al. (2021); using a MAS system on a Bayesian network test-bed to distinguish between truthful and deceptive agents based on the correlation of the agents’ beliefs Santos and Li (2009); modelling agent based deceptive interactions on social networks Barrio et al. (2015); modelling deceptive interactions to counter reconnaissance-based cyber-attacks Schlenker et al. (2018); studying the effects of deception in repeated games between learning agents Nguyen et al. (2019); applying IDT in multi-agent reinforcement learning to reduce the effects of social engineering attacks Yang and Yu (2018); developing an information-theoretic model of deceptive strategies Kopp et al. (2018); learning deceptive strategies through bayesian belief manipulation Aitchison et al. (2021); formalising cyber-deception games between multiple agents using hyper-game theory methods Ferguson-Walter et al. (2019). Additionally, given the increasing interest in explainability for AI, Wright et al. (2019) discuss how agents can strategically provide users with deceptive and rebellious explanations.
@@ 3.2.4 Reasoning in Agent Deception
Work in this area of deceptive AI has explored issues such as using and detecting deception in argument debate games by Sakama (2012) that are formalised using abstract argumentation Dung (1995); using abductive reasoning for deception Sakama (2011); using argument mining for detecting deceptive reviews online Cocarascu and Toni (2016); using argument mining for detecting propaganda Vorakitphan et al. (2021); using argumentation-based tools to analyse disinformation in fake news Delobelle et al. (2020); looking into what type of arguments can be used by a machine to deceive Clark (2010); modelling agents that use mindreading for deception Isaac and Bridewell (2014); modelling the meta-reasoning of agents that tell deceptive stories and detect deceptive stories in dialogue argumentation games Sarkadi et al. (2019a).
@@ 3.2.5 Engineering Deceptive Agents
Recent work in this area has looked at issues such as: modelling agent architectures that engage in deceptive storytelling during dialogues Rato et al. (2017); developing deceptive chatbots for emotional well-being and mental health Galindo et al. (2021); engineering agents that use lies, bullshit and deception Panisson et al. (2018a); and also the first attempt in MAS to integrate IMT2, TDT, and IDT in order to model, implement and evaluate deceptive interactions between complex reasoning agents that use Theory of Mind by Sarkadi et al. (2019b).
@@ 3.2.6 Embodied Deceptive Agents
Research in this area looks at issues such as: how deception about agency influences the way children treat robots on a social level and when is it acceptable to deceive children about the nature of the robots they interact with Westlund and Breazeal (2015); how robots can decide to perform physical actions such that they manipulate the beliefs of humans who observe them Gray and Breazeal (2014); how to provide robots with the capacity to deceive Wagner and Arkin (2011); why the properties of deception and transparency in robots should be treated in the context of user vulnerability Collins (2017); using biologically inspired behaviour for robot deception Shim and Arkin (2012), such as squirrel behaviour and applying it to shill agents Pettinati et al. (2021). Shim and Arkin (2013) even provide a taxonomy of robot deception and present the related benefits of deception for human-robot interaction.
@@
We have seen in this section that research on deception involves looking at various aspects and components of deception. However, more often than not, AI does so from a highly reductionist approach. This type of reductionism can be problematic, because compared to other forms of dishonesty such as lying or bullshitting, both of which can be elegantly modelled in logics of communication, deception involves more complex cognitive mechanisms that have to be modelled Panisson et al. (2018a); Hyman (1990), usually in a relational manner McHugh et al. (2004). For instance, ToM is considered a critical component of deception Sarkadi et al. (2019b); Sarkadi (2021); Isaac and Bridewell (2017); De Rosis et al. (2003), and more often than not, models in AI fail to consider ToM as a complex form of socio-cognitive process [^4] The AI literature has also emphasised the potential benefits of artificial agents with ToM (that is not necessarily high-level simulation), that include the explainability, the efficiency, and the increased social performance of machines. For example, de Weerd et al. (2017) show how agents with higher order ToM (‘higher’ here means strictly meta-level recursivity over a single belief, not in the sense of ‘high-level simulation’) outperform other agents in negotiation. Another ability is increasing the seamlessness of cooperation between humans and robots Winfield (2018); Bianco and Ognibene (2020). Such benefits, we believe, might be the reason why considerable efforts are being made in the AI community to enable machines to form and use models of other minds Albrecht and Stone (2018)..
@@
In hybrid societies, modelling the ToM component has both risks and benefits, as it allows us to give agents the ability to align their ethical values with humans Isaac and Bridewell (2017), but then it also allows machines to apply higher levels of ToM to outperform humans in deception and counter-deception Sarkadi (2021). What we mean here by higher level is either deeper levels of recursivity or a greater number of parallel simulation instances/branches when modelling other minds in the sense of Goldman’s ‘high-level simulation’. As an example, machines could outperform humans using higher-level ToM by exploiting unknown unknowns Sarkadi et al. (2019b) [^5] Reports that say that something hasn’t happened are always interesting to me, because as we know, there are known knowns; there are things we know we know. We also know there are known unknowns; that is to say we know there are some things we do not know. But there are also unknown unknowns—the ones we don’t know we don’t know. And if one looks throughout the history of our country and other free countries, it is the latter category that tends to be the difficult ones.Rumsfeld (2002).. Introduced publicly during a US Dept. of Defense news briefing by Donald Rumsfeld in the context of Intelligence Analysis, the unknown unknown means not knowing that one does not know something Logan (2009); Rumsfeld (2002). By exploiting unknown unknowns, AI agents do not simply exploit missing information due to the target’s inaccessibility to sources, but also the target’s failure to reflect on the information that is or is not available - either due to bias or due to limited cognitive capabilities or physical resources (lack of time or lack of analysis tools). Conclusively, it is unrealistic to model deception in agent societies without having the proper kind of representation and implementation for the deceptive agents, their internal cognitive processes, and their behaviour within the complex system in which they interact.
@@
Agent Societies Mell et al. (2018); Ishowo-Oloko et al. (2019); Dras et al. (2010); Sarkadi (2024)
Logical Aspects Sakama et al. (2010); Sakama (2015); Sakama et al. (2015); Sakama (2021); Bonnet et al. (2021); Smith et al. (2016); Van Ditmarsch (2014); Uckelman (2011); Jones (2015)
Agent Strategies Christian and Young (2004); Masters and Sardina (2017); Aitchison et al. (2021); Benke et al. (2021); Santos and Li (2009); Barrio et al. (2015); Schlenker et al. (2018); Nguyen et al. (2019); Yang and Yu (2018); Kopp et al. (2018); Ferguson-Walter et al. (2019); Wright et al. (2019)
Reasoning Sakama (2012); Dung (1995); Sakama (2011); Cocarascu and Toni (2016); Clark (2010); Isaac and Bridewell (2014); Sarkadi et al. (2019a)
Engineering Rato et al. (2017); Galindo et al. (2021); Panisson et al. (2018a); Sarkadi et al. (2019b)
Embodiment Westlund and Breazeal (2015); Gray and Breazeal (2014); Wagner and Arkin (2011); Shim and Arkin (2012, 2013); Pettinati et al. (2021)
Ethics* Kampik et al. (2018); Sklar et al. (2004); Isaac and Bridewell (2017); Collins (2017); Chakraborti and Kambhampati (2019); Coeckelbergh (2018); Shim and Arkin (2013); Sharkey and Sharkey (2021); Natale et al. (2021); Bender et al. (2021); Sarkadi et al. (2023); Sarkadi (2023a)
Table 1: The research categories on modelling deception with AI techniques described in section 3. For a holistic overview of ethics and Deceptive AI, see Sarkadi (2023a).
@@ 4 DAMAS: A MAS Framework for Deception Analysis
First, such a framework should offer enough representational power in order to enable the automation of counterfactual reasoning about social dynamics between agents. This is achieved by representing (i) the environment, (ii) agents and their communicative abilities, (iii) the knowledge that agents have (such as nested beliefs and relations between beliefs and ToM of each other), (iv) how changes in the environment influence the agents’ beliefs, and (v) how the agents themselves cause these changes intentionally or unintentionally - all of these in a relational fashion McHugh et al. (2004).
Second, the framework should be both scalable and modular. In other words, it needs to be able to support the seamless addition and integration of new agent, environment, and organisation types. That is, the models of the world should be easily increased and modified based on potential changes in deception modelling requirements. For instance, whenever a new deception model or a new type of deceptive agent architecture is developed, then this should be easily added to the framework without changing the already existing models or agent architectures. The interoperability of agents and tools can guarantee this property. MAS frameworks such as JaCaMo are designed based on these two principles, enabling an almost seamless interoperability between the environment, agent, and organisation layers Boissier et al. (2013).
Third, such a framework should be usable. Analysts should be able to use the framework to explain why or why not interactions between agents happen the way they happen. Explainable AI (XAI) techniques can be used to extract narratives of agent dynamics and generate explanations in a human-interpretable manner, e.g. by taking into account the principles of causality, counterfactuality, contrastiveness, and social factors to tailor the explanations of agent behaviour based on the user’s background knowledge, expertise, and preferences Miller (2019).
@@ 4.1.1 Interaction (MAS) Models for DAMAS
A strong focus in AI is on the logical formalisation and categorisation of one-way deception. For example, Sakama and Caminada (2010) aim to categorise multiple forms of one-way deception and explain their logical properties, and Sakama et al. (2010) look at the multiple variations of dishonesty including one-way deception, whereas Caminada (2009) looks at the logical relation between truth and other forms of dishonesty such as lying, bullshitting [^6] Bullshitting, according to Frankfurt (2009) is making statements without regard for their truth value. and even some forms of one-way deception. Sakama (2015) also uses dynamic epistemic logic to formalise deceptive acts of communication. Later, Sakama (2021) adds a causal model for epistemic dynamics determined by communication.
One-way deception has been also studied from an agent-oriented perspective. De Rosis et al. (2003) show a simulation of deception in belief networks, whereas Christian and Young (2004) define a model of a heuristic plan search for finding a deception plan. More recently, Panisson et al. (2018a) defined and implemented a BDI agent using Jason (an agent oriented programming language) that can choose to lie, bullshit or deceive in order to manipulate the beliefs of another agent. This work was continued by Sarkadi et al. (2019b); Sarkadi (2021) who integrated it with TDT, IDT and IMT2. There, the deceptive interactions between two BDI agents are defined and implemented, where the deceiver agent simulates the mind of its target taking into account the levels of the target’s trust, the confidence in its ToM of the target, and the level of its communicative skill. The approach introduces the idea that deception can be analysed in multi-agent social interactions not only regarding its success (successful vs unsuccessful), but also regarding whether it was attempted or intended (intended vs unintended).
@@
Work in the Intelligence Analysis literature has led to solid psychological theories of counter-deception and deception detection indicating that Intelligence and Espionage Agencies are often engaging in very complex counter-deception analysis Heuer (1999). Counter deception has also found its applications in interrogations. When interrogators happen to deal with deceptive or manipulative subjects, they can resort to counter-deception to increase their chances at successful interrogation Vrij and Granhag (2012). For example, the interrogator can pretend to know some information a priori such that the subject is tricked into giving answers that reveal or confirm the truth of the information to the interrogator’s questions. Studies show that interrogators trained in counter-deception have a greater success at deception-detection Hartwig et al. (2006). Moreover, this effect can be also observed in MAS simulations, where deceivers and interrogators seem to engage in a ToM arms race Sarkadi (2023b).
@@
- In theory, counter deception can be infinitely recursive. The property of recursivity in counter-deception, however, depends on the type of ToM of the opponent. A recursive ToM means that agents add levels of ToM on top of each other’s ToM of themselves. For example, “I know that you know that I know…ad infinitum…some information” represents a recursive ToM. An entirely recursive ToM would mean that the deceptive reasoning processes of a deceiver’s mind would only focus on taking a certain belief, let us say $Bel_{i}(P)$ and infering $Bel_{j}^{k}(Bel_{i}^{k}(P))$ in order to gain some advantage using deception, where $i$ and $j$ represent different agents and $k$ represents the level of ToM. However, in interactions that assume an entirely recursive ToM, the only advantage belongs to the agent that has a the greater level of ToM as shown by agent-based simulations of games played between agents with multiple orders of ToM de Weerd et al. (2017).
@@
- In practice, human agents are rationally bounded and are not capable of infinitely recursive reasoning. Thus, in real life, deception is not applied to infinitely recursive mental models. There might be cases in which a deceiver could exploit its target’s mind without engaging in expensive recursive reasoning. There might be beliefs that do not exist in the target’s ToM of the deceiver. Or there might be beliefs of the deceiver that the target does not know that it (the target) does not know (as in the case of Rumsfeld’s unknown unknowns). In such cases, it might be wiser for the deceiver to avoid spending cognitive resources on the higher-order reasoning of ToM and exploit other types of beliefs inside the target’s mind. The deceiver might, for example, simulate the belief updates that happen in the mind of the target in order to see what new beliefs can be formed and also explore which of these newly formed beliefs can be more efficiently exploited. This type of ToM implies a dynamic semantic ToM model. Dynamic semantic models of ToM in MAS based on belief-desire-intention architectures and agent oriented communication along with their use under uncertainty have been introduced and implemented by Sarkadi et al. (2018), and more recently described by Shvo et al. (2020). One instance of partially recursive counter-deception has also been introduced as preliminary work on argumentation dialogue games, where both the deceiver and the interrogator have a ToM of each other that they update after every interaction, in Sarkadi et al. (2019a). The deceiver uses its ToM to build a story that forces the interrogator to accept it, and viceversa, the interrogator forces the deceiver to accept that it has not found a believable story.
@@
- The exception to the presumably intuitive rule that deception requires at least two agents (deceiver and target) is the case of self-deception. In order for self-deception to be successful, the deceiver must be able to deceive itself, playing both the role of the deceiver and its target. Here we have a paradoxical situation. Assuming that the deceiver needs a ToM of its target in order to deceive, then the deceiver needs a ToM of itself. Given that the same entity plays both the roles, then it must ontologically follow that its ToM of itself must be complete, i.e. there is no knowledge about itself that it does not know. If the ToM is complete, then the ToM must include the deceiver’s deceptive intentions or goals as well as the target’s goal of not being deceived. Obviously, these two types of conflicting goals and intentions determine an inconsistent system. However, there are some special cases in which these paradoxical situations can be overcome. Jones (2015) manages to model a specific set of cases of self-deception that are logically consistent in Hintikka’s logic of belief. The author does emphasise that such inconsistencies still remain inside the system, but become latent due to the specific cases that are formalised, and Jones explains the reasons for them becoming latent. A fairly new idea has been proposed in by Chen (2021) to use models of self-deception in designing persuasive software applications to nudge people to sleep better, but even though it is worth mentioning, this type of research is still in its infancy. On a similar note, before Chen, Borenstein and Arkin (2016) have actually proposed the idea that ethical robots could be used to nudge humans into making more ethical decisions in their life. Perhaps self-deception can actually drive teamwork and cooperation in hybrid societies? If so, then how do we model interactions to analyse the outcomes of promoting self-deception in humans with AI technologies, or even deploying self-deceptive agents?

### v2

@@ 3.2 Previous Work
In the last 20 years, the importance of research on deceptive AI has been emphasised in the AI community during a series of events, namely the 2015 AAAI Fall Symposium on Deceptive and Counter-Deceptive Machines [^2] See the proceedings at https://aaai.org/Press/Reports/Symposia/Fall/fs-15-03.php., and the two International Workshops on Deceptive AI [^3] See the joint proceedings in (Sarkadi et al., 2021b). co-located with ECAI2020 and IJCAI2021. Having gone through the socio-cognitive components that need to be considered when modelling deceptive AI, we now describe the work at the intersection of deception and AI, grouping them into six categories to better understand what computational aspects and properties deception they address.
@@ 3.2.1 Deception in Agent Societies
Recent findings back up this perspective by showing how humans and artificial agents influence each other’s attitudes and behaviour. For example, (Mell et al., 2018) focus on how the negotiation strategies of artificial agents determine humans to endorse artificial deception. Apparently and counter-intuitively, humans also tend to be more cooperative with deceptive machines than with friendly ones, according to (Ishowo-Oloko et al., 2019). Another type of study, presented by (Dras et al., 2010), looks at the ways in which deceptive language can be generated and detected in virtual systems.
@@ 3.2.2 Logical Aspects of Agent Deception
Works in the areas of Logic and Philosophy have strongly influenced the current state of the art approaches in deceptive AI. Most known are the works in knowledge representation, which have laid the foundations to formalise deceptive communication. These appear in (Sakama et al., 2010), in (Sakama, 2015), and in (Sakama et al., 2015), where the authors present the logical distinctions between different types of dishonest linguistic behaviour, and later introducing a model of causality for them (Sakama, 2021);(Bonnet et al., 2021) sketch a logical theory of belief manipulation; (Van Ditmarsch, 2014) introduce a dynamic logic of lying that considers lies about factual propositions as well as lies about the beliefs of others; in (Uckelman, 2011) where the author uses the medieval concept of dubitatio to study deceptive agents; (Jones, 2015) proposes a formal model of self-deception that is consistent given certain epistemic conditions imposed on the agent that engages in self-deception; and (Smith et al., 2016) describe a logic-based account of analysing conjuring tricks as form of deception, and identify substitutable elements and stable occlusion as crucial components to construct perceived impossibilities.
@@ 3.2.3 Strategies for Agent Deception
Some of the work in this area explores issues such as using heuristics to cause target agents to execute a plan that will achieve the deceiver’s desired goal (Christian and Young, 2004); finding deceptive strategies using path-planning (Masters and Sardina, 2017; Masters et al., 2021a; Benke et al., 2021); using a MAS system on a Bayesian network test-bed to distinguish between truthful and deceptive agents based on the correlation of the agents’ beliefs (Santos and Li, 2009); modelling agent based deceptive interactions on social networks (Barrio et al., 2015); modelling deceptive interactions to counter reconnaissance-based cyber-attacks (Schlenker et al., 2018); studying the effects of deception in repeated games between learning agents (Nguyen et al., 2019); applying IDT in multi-agent reinforcement learning to reduce the effects of social engineering attacks (Yang and Yu, 2018); developing an information-theoretic model of deceptive strategies (Kopp et al., 2018); learning deceptive strategies through bayesian belief manipulation (Aitchison et al., 2021); formalising cyber-deception games between multiple agents using hyper-game theory methods (Ferguson-Walter et al., 2019). Additionally, given the increasing interest in explainability for AI, (Wright et al., 2019) discuss how agents can strategically provide users with deceptive and rebellious explanations.
@@ 3.2.4 Reasoning in Agent Deception
Work in this area of deceptive AI has explored issues such as using and detecting deception in argument debate games by (Sakama, 2012) that are formalised using abstract argumentation (Dung, 1995); using abductive reasoning for deception (Sakama, 2011); using argument mining for detecting deceptive reviews online (Cocarascu and Toni, 2016); using argument mining for detecting propaganda (Vorakitphan et al., 2021); using argumentation-based tools to analyse disinformation in fake news (Delobelle et al., 2020); looking into what type of arguments can be used by a machine to deceive (Clark, 2010); modelling agents that use mindreading for deception (Isaac and Bridewell, 2014); modelling the meta-reasoning of agents that tell deceptive stories and detect deceptive stories in dialogue argumentation games (Sarkadi et al., 2019a).
@@ 3.2.5 Engineering Deceptive Agents
Recent work in this area has looked at issues such as: modelling agent architectures that engage in deceptive storytelling during dialogues (Rato et al., 2017); developing deceptive chatbots for emotional well-being and mental health (Galindo et al., 2021); engineering agents that use lies, bullshit and deception (Panisson et al., 2018a); and also the first attempt in MAS to integrate IMT2, TDT, and IDT in order to model, implement and evaluate deceptive interactions between complex reasoning agents that use Theory of Mind by (Sarkadi et al., 2019b).
@@ 3.2.6 Embodied Deceptive Agents
Research in this area looks at issues such as: how deception about agency influences the way children treat robots on a social level and when is it acceptable to deceive children about the nature of the robots they interact with (Westlund and Breazeal, 2015); how robots can decide to perform physical actions such that they manipulate the beliefs of humans who observe them (Gray and Breazeal, 2014); how to provide robots with the capacity to deceive (Wagner and Arkin, 2011); why the properties of deception and transparency in robots should be treated in the context of user vulnerability (Collins, 2017); using biologically inspired behaviour for robot deception (Shim and Arkin, 2012), such as squirrel behaviour and applying it to shill agents (Pettinati et al., 2021). (Shim and Arkin, 2013) even provide a taxonomy of robot deception and present the related benefits of deception for human-robot interaction.
@@
We have seen in this section that research on deception involves looking at various aspects and components of deception. However, more often than not, AI does so from a highly reductionist approach. This type of reductionism can be problematic, because compared to other forms of dishonesty such as lying or bullshitting, both of which can be elegantly modelled in logics of communication, deception involves more complex cognitive mechanisms that have to be modelled (Panisson et al., 2018a; Hyman, 1990), usually in a relational manner (McHugh et al., 2004). For instance, ToM is considered a critical component of deception (Sarkadi et al., 2019b; Sarkadi, 2021; Isaac and Bridewell, 2017; De Rosis et al., 2003), and more often than not, models in AI fail to consider ToM as a complex form of socio-cognitive process [^4] The AI literature has also emphasised the potential benefits of artificial agents with ToM (that is not necessarily high-level simulation), that include the explainability, the efficiency, and the increased social performance of machines. For example, (de Weerd et al., 2017) show how agents with higher order ToM (‘higher’ here means strictly meta-level recursivity over a single belief, not in the sense of ‘high-level simulation’) outperform other agents in negotiation. Another ability is increasing the seamlessness of cooperation between humans and robots (Winfield, 2018; Bianco and Ognibene, 2020). Such benefits, we believe, might be the reason why considerable efforts are being made in the AI community to enable machines to form and use models of other minds (Albrecht and Stone, 2018)..
@@
In hybrid societies, modelling the ToM component has both risks and benefits, as it allows us to give agents the ability to align their ethical values with humans (Isaac and Bridewell, 2017), but then it also allows machines to apply higher levels of ToM to outperform humans in deception and counter-deception (Sarkadi, 2021). What we mean here by higher level is either deeper levels of recursivity or a greater number of parallel simulation instances/branches when modelling other minds in the sense of Goldman’s ‘high-level simulation’. As an example, machines could outperform humans using higher-level ToM by exploiting unknown unknowns (Sarkadi et al., 2019b) [^5] Reports that say that something hasn’t happened are always interesting to me, because as we know, there are known knowns; there are things we know we know. We also know there are known unknowns; that is to say we know there are some things we do not know. But there are also unknown unknowns—the ones we don’t know we don’t know. And if one looks throughout the history of our country and other free countries, it is the latter category that tends to be the difficult ones.(Rumsfeld, 2002).. Introduced publicly during a US Dept. of Defense news briefing by Donald Rumsfeld in the context of Intelligence Analysis, the unknown unknown means not knowing that one does not know something (Logan, 2009; Rumsfeld, 2002). By exploiting unknown unknowns, AI agents do not simply exploit missing information due to the target’s inaccessibility to sources, but also the target’s failure to reflect on the information that is or is not available - either due to bias or due to limited cognitive capabilities or physical resources (lack of time or lack of analysis tools). Conclusively, it is unrealistic to model deception in agent societies without having the proper kind of representation and implementation for the deceptive agents, their internal cognitive processes, and their behaviour within the complex system in which they interact.
@@
Agent Societies (Mell et al., 2018; Ishowo-Oloko et al., 2019; Dras et al., 2010; Sarkadi, 2024)
Logical Aspects (Sakama et al., 2010; Sakama, 2015; Sakama et al., 2015; Sakama, 2021; Bonnet et al., 2021; Smith et al., 2016; Van Ditmarsch, 2014; Uckelman, 2011; Jones, 2015)
Agent Strategies (Christian and Young, 2004; Masters and Sardina, 2017; Aitchison et al., 2021; Benke et al., 2021; Santos and Li, 2009; Barrio et al., 2015; Schlenker et al., 2018; Nguyen et al., 2019; Yang and Yu, 2018; Kopp et al., 2018; Ferguson-Walter et al., 2019; Wright et al., 2019)
Reasoning (Sakama, 2012; Dung, 1995; Sakama, 2011; Cocarascu and Toni, 2016; Clark, 2010; Isaac and Bridewell, 2014; Sarkadi et al., 2019a)
Engineering (Rato et al., 2017; Galindo et al., 2021; Panisson et al., 2018a; Sarkadi et al., 2019b)
Embodiment (Westlund and Breazeal, 2015; Gray and Breazeal, 2014; Wagner and Arkin, 2011; Shim and Arkin, 2012, 2013; Pettinati et al., 2021)
Ethics* (Kampik et al., 2018; Sklar et al., 2004; Isaac and Bridewell, 2017; Collins, 2017; Chakraborti and Kambhampati, 2019; Coeckelbergh, 2018; Shim and Arkin, 2013; Sharkey and Sharkey, 2021; Natale et al., 2021; Bender et al., 2021; Sarkadi et al., 2023; Sarkadi, 2023a)
Table 1: The research categories on modelling deception with AI techniques described in section 3. For a holistic overview of ethics and Deceptive AI, see (Sarkadi, 2023a).
@@ 4 DAMAS: A MAS Framework for Deception Analysis
First, such a framework should offer enough representational power in order to enable the automation of counterfactual reasoning about social dynamics between agents. This is achieved by representing (i) the environment, (ii) agents and their communicative abilities, (iii) the knowledge that agents have (such as nested beliefs and relations between beliefs and ToM of each other), (iv) how changes in the environment influence the agents’ beliefs, and (v) how the agents themselves cause these changes intentionally or unintentionally - all of these in a relational fashion (McHugh et al., 2004).
Second, the framework should be both scalable and modular. In other words, it needs to be able to support the seamless addition and integration of new agent, environment, and organisation types. That is, the models of the world should be easily increased and modified based on potential changes in deception modelling requirements. For instance, whenever a new deception model or a new type of deceptive agent architecture is developed, then this should be easily added to the framework without changing the already existing models or agent architectures. The interoperability of agents and tools can guarantee this property. MAS frameworks such as JaCaMo are designed based on these two principles, enabling an almost seamless interoperability between the environment, agent, and organisation layers (Boissier et al., 2013).
Third, such a framework should be usable. Analysts should be able to use the framework to explain why or why not interactions between agents happen the way they happen. Explainable AI (XAI) techniques can be used to extract narratives of agent dynamics and generate explanations in a human-interpretable manner, e.g. by taking into account the principles of causality, counterfactuality, contrastiveness, and social factors to tailor the explanations of agent behaviour based on the user’s background knowledge, expertise, and preferences (Miller, 2019).
@@ 4.1.1 Interaction (MAS) Models for DAMAS
A strong focus in AI is on the logical formalisation and categorisation of one-way deception. For example, (Sakama and Caminada, 2010) aim to categorise multiple forms of one-way deception and explain their logical properties, and (Sakama et al., 2010) look at the multiple variations of dishonesty including one-way deception, whereas (Caminada, 2009) looks at the logical relation between truth and other forms of dishonesty such as lying, bullshitting [^6] Bullshitting, according to (Frankfurt, 2009) is making statements without regard for their truth value. and even some forms of one-way deception. (Sakama, 2015) also uses dynamic epistemic logic to formalise deceptive acts of communication. Later, (Sakama, 2021) adds a causal model for epistemic dynamics determined by communication.
One-way deception has been also studied from an agent-oriented perspective. (De Rosis et al., 2003) show a simulation of deception in belief networks, whereas (Christian and Young, 2004) define a model of a heuristic plan search for finding a deception plan. More recently, (Panisson et al., 2018a) defined and implemented a BDI agent using Jason (an agent oriented programming language) that can choose to lie, bullshit or deceive in order to manipulate the beliefs of another agent. This work was continued by (Sarkadi et al., 2019b; Sarkadi, 2021) who integrated it with TDT, IDT and IMT2. There, the deceptive interactions between two BDI agents are defined and implemented, where the deceiver agent simulates the mind of its target taking into account the levels of the target’s trust, the confidence in its ToM of the target, and the level of its communicative skill. The approach introduces the idea that deception can be analysed in multi-agent social interactions not only regarding its success (successful vs unsuccessful), but also regarding whether it was attempted or intended (intended vs unintended).
@@
Work in the Intelligence Analysis literature has led to solid psychological theories of counter-deception and deception detection indicating that Intelligence and Espionage Agencies are often engaging in very complex counter-deception analysis (Heuer, 1999). Counter deception has also found its applications in interrogations. When interrogators happen to deal with deceptive or manipulative subjects, they can resort to counter-deception to increase their chances at successful interrogation (Vrij and Granhag, 2012). For example, the interrogator can pretend to know some information a priori such that the subject is tricked into giving answers that reveal or confirm the truth of the information to the interrogator’s questions. Studies show that interrogators trained in counter-deception have a greater success at deception-detection (Hartwig et al., 2006). Moreover, this effect can be also observed in MAS simulations, where deceivers and interrogators seem to engage in a ToM arms race (Sarkadi, 2023b).
@@
- In theory, counter deception can be infinitely recursive. The property of recursivity in counter-deception, however, depends on the type of ToM of the opponent. A recursive ToM means that agents add levels of ToM on top of each other’s ToM of themselves. For example, “I know that you know that I know…ad infinitum…some information” represents a recursive ToM. An entirely recursive ToM would mean that the deceptive reasoning processes of a deceiver’s mind would only focus on taking a certain belief, let us say $Bel_{i}(P)$ and infering $Bel_{j}^{k}(Bel_{i}^{k}(P))$ in order to gain some advantage using deception, where $i$ and $j$ represent different agents and $k$ represents the level of ToM. However, in interactions that assume an entirely recursive ToM, the only advantage belongs to the agent that has a the greater level of ToM as shown by agent-based simulations of games played between agents with multiple orders of ToM (de Weerd et al., 2017).
@@
- In practice, human agents are rationally bounded and are not capable of infinitely recursive reasoning. Thus, in real life, deception is not applied to infinitely recursive mental models. There might be cases in which a deceiver could exploit its target’s mind without engaging in expensive recursive reasoning. There might be beliefs that do not exist in the target’s ToM of the deceiver. Or there might be beliefs of the deceiver that the target does not know that it (the target) does not know (as in the case of Rumsfeld’s unknown unknowns). In such cases, it might be wiser for the deceiver to avoid spending cognitive resources on the higher-order reasoning of ToM and exploit other types of beliefs inside the target’s mind. The deceiver might, for example, simulate the belief updates that happen in the mind of the target in order to see what new beliefs can be formed and also explore which of these newly formed beliefs can be more efficiently exploited. This type of ToM implies a dynamic semantic ToM model. Dynamic semantic models of ToM in MAS based on belief-desire-intention architectures and agent oriented communication along with their use under uncertainty have been introduced and implemented by (Sarkadi et al., 2018), and more recently described by (Shvo et al., 2020). One instance of partially recursive counter-deception has also been introduced as preliminary work on argumentation dialogue games, where both the deceiver and the interrogator have a ToM of each other that they update after every interaction, in (Sarkadi et al., 2019a). The deceiver uses its ToM to build a story that forces the interrogator to accept it, and viceversa, the interrogator forces the deceiver to accept that it has not found a believable story.
@@
- The exception to the presumably intuitive rule that deception requires at least two agents (deceiver and target) is the case of self-deception. In order for self-deception to be successful, the deceiver must be able to deceive itself, playing both the role of the deceiver and its target. Here we have a paradoxical situation. Assuming that the deceiver needs a ToM of its target in order to deceive, then the deceiver needs a ToM of itself. Given that the same entity plays both the roles, then it must ontologically follow that its ToM of itself must be complete, i.e. there is no knowledge about itself that it does not know. If the ToM is complete, then the ToM must include the deceiver’s deceptive intentions or goals as well as the target’s goal of not being deceived. Obviously, these two types of conflicting goals and intentions determine an inconsistent system. However, there are some special cases in which these paradoxical situations can be overcome. (Jones, 2015) manages to model a specific set of cases of self-deception that are logically consistent in Hintikka’s logic of belief. The author does emphasise that such inconsistencies still remain inside the system, but become latent due to the specific cases that are formalised, and Jones explains the reasons for them becoming latent. A fairly new idea has been proposed in by (Chen, 2021) to use models of self-deception in designing persuasive software applications to nudge people to sleep better, but even though it is worth mentioning, this type of research is still in its infancy. On a similar note, before Chen, (Borenstein and Arkin, 2016) have actually proposed the idea that ethical robots could be used to nudge humans into making more ethical decisions in their life. Perhaps self-deception can actually drive teamwork and cooperation in hybrid societies? If so, then how do we model interactions to analyse the outcomes of promoting self-deception in humans with AI technologies, or even deploying self-deceptive agents?